Consecutive even integers find four consecutive even intergers whose sum is 340
Answer by checkley77(12844)

X, X+2, X+4, X+6 ARE THE INTEGERS.
X+X+2+X+4+X+6=340
4X+12=340
4X=340-12
4X=328
X=328/4
X=82 FOR THE FIRST INTEGER.
84,86 & 88 ARE THE REST OF THE INTEGERS.
PROOF:
82+84+86+88=340
340=340